# Dependency-Aware Semi-Structured Sparsity of GLU Variants in Large Language Models

### v1

@@ Abstract
The rapid advancement in Large Language Models (LLMs) has markedly enhanced the capabilities of language understanding and generation. However, the substantial model size poses hardware challenges, affecting both memory size for serving and inference latency for token generation. To address those challenges, we propose Dependency-aware Semi-structured Sparsity (DaSS), a novel method for the recent prevalent SwiGLU-based LLMs pruning. Our approach incorporates structural dependency into the weight magnitude-based unstructured pruning. We introduce an MLP-specific pruning metric that evaluates the importance of each weight by jointly considering its magnitude and its corresponding MLP intermediate activation norms. DaSS facilitates a balance between the adaptability offered by unstructured pruning and the structural consistency inherent in dependency-based structured pruning. Empirical evaluations on Mistral and LLaMA2 model families demonstrate that DaSS not only outperforms both SparseGPT and Wanda in achieving hardware-friendly N:M sparsity patterns but also maintains the computational efficiency of Wanda.
Machine Learning, ICML
@@ 1 Introduction
Recent years have witnessed the great success of Transformer-based Large Language Models (LLMs) (Brown et al., 2020; Chowdhery et al., 2023; OpenAI, 2023) across various challenging tasks, such as mathematical reasoning (Cobbe et al., 2021), code generation (Chen et al., 2021). However, the practical use of these models for inference has faced a major obstacle due to the substantial computational resources they consume. To tackle this, several methods for efficient LLMs inference have been applied, including training compact models using more data (Touvron et al., 2023a), model quantization (Dettmers et al., 2022; Frantar et al., 2023), and grouped query attention (Shazeer, 2019; Ainslie et al., 2023).
Neural network pruning (Han et al., 2015; Wen et al., 2016) is a widely adopted technique for compressing models, leading to a substantial reduction in model size and inference latency. Pruning can be categorized into two main approaches: unstructured pruning (Sun et al., 2023; Frantar & Alistarh, 2023), which involves the removal of specific weights, and structured pruning (Zhang et al., 2023; Ma et al., 2023), which entails the removal of complete rows or columns of weights. In contrast to structured pruning, which struggles with performance in LLMs even at low sparsity levels, unstructured pruning (Sun et al., 2023; Frantar & Alistarh, 2023) exhibits promising results without additional retraining, and achieves practical speedup in Nvidia GPU through more stringent N:M sparsity pattern (Mishra et al., 2021; Kurtic et al., 2023). SparseGPT (Frantar & Alistarh, 2023) tackles the issue of pruning LLMs by considering it as a problem of layerwise reconstruction. It exclusively depends on weight updates aimed at maintaining the input-output relationship for each layer. Motivated by the emergent large magnitude activations in LLMs (Dettmers et al., 2022), Wanda (Sun et al., 2023) further eases the computational load in SparseGPT by solely employing the product derived from weights and input activation magnitudes.
@@
Figure 1: Illustration of our proposed Dependency-aware Semi-structured Sparsity (DaSS). In (a) dependency-based structured pruning (Ma et al., 2023), all the weights connecting to the same intermediate neuron are removed or remain simultaneously. In (b) Wanda unstructured pruning (Sun et al., 2023), it assigns greater emphasis to the weights corresponding to large input activations. For Gate-Proj and UP-Proj, the same number of weights are removed for each MLP neuron, regardless of whether some neurons have much larger activation norms. For Down-Proj, the weights corresponding to larger activation norms are more likely to be pruned. This can lead to a structural mismatch. In (c), all the weights corresponding to large intermediate activations are more likely to be reserved.
As the SwiGLU-based MLP (Shazeer, 2020) module accounts for more than 80% parameters in the recent LLMs that use multi/grouped query attention (Chowdhery et al., 2023; Touvron et al., 2023b) [^1] In LLaMA2-70B, the dimension of key and value is $\frac{1}{8}d$, MLP intermediate dimension is $\frac{7}{2}d$., its pruning emerges as a pivotal factor in determining the overall compression efficacy of LLMs. In dependency-aware structured pruning, it’s critical to consider that pruned parameters have dependencies with other parameters, owing to their interconnected nature (Ma et al., 2023; Fang et al., 2023). In the context of MLP pruning, all the weights connected to each intermediate neuron should be preserved or pruned simultaneously. The precise coordination involved in pruning is crucial for upholding the model’s structural integrity and its functional capabilities. Although current unstructured pruning methods (Sun et al., 2023; Frantar & Alistarh, 2023) effectively remove a significant number of redundant weights, they operate entirely locally within each linear layer without considering inter-dependencies to other layers. This can lead to a structural mismatch, which is more evident in Wanda as shown in Figure 1(b): In Gate and Up projections, the same amount of parameters are pruned for each MLP neuron. However, the intermediate activation norms of SwiGLU are not uniformly distributed and some neurons have much larger norms than others. Based on Wanda pruning metric, more weights connected to neurons with large activation norms are preserved. At high sparsity, this problem gets magnified in Wanda causing significant drops in performance by the broken network.
@@
In order to overcome the limitations present in current pruning methodologies, we introduce a new paradigm, namely, Dependency-aware Semi-structured Sparsity (DaSS). This approach is specifically designed to navigate the middle ground between the flexibility of unstructured pruning and the structural consistency of dependency-based structured pruning. We examine the effectiveness of magnitude pruning in Gate and Up projections of SwiGLU, and we find that grouping weights per-input is surprisingly effective compared with Wanda (Sun et al., 2023). Leveraging these newly acquired understandings, we present a new MLP pruning metric that assesses each weight’s importance based on the product of its magnitude and the norm of the corresponding MLP intermediate activations. Our proposed DaSS method, illustrated in Figure 1(c), embodies a semi-structured pattern that retains a degree of the adaptability inherent in unstructured pruning while incorporating the dependency-aware aspect of structured pruning. This balance allows for more precise pruning. DaSS can be easily extended to hardware-friendly N:M sparsity patterns.
We perform extensive experiments on LLaMA2 and Mistral to evaluate DaSS across various tasks from language modeling, commonsense reasoning, and MMLU (Hendrycks et al., 2021). In achieving hardware-friendly N:M sparsity patterns, DaSS consistently excels beyond the existing LLM pruning methods SparseGPT (Frantar & Alistarh, 2023) and Wanda (Sun et al., 2023), while maintaining the computational efficiency akin to Wanda. Impressively, DaSS outperforms SparseGPT at high sparsity even without weight update. Our work contributes fresh insights into the nuanced understanding of sparsity within large language models.
@@ 2.1 Wanda Pruning Method
In the context of LLM pruning, we denote a linear layer weight matrix $\mathbf{W}\in\mathbb{R}^{d_{\text{out}}\times d_{\text{in}}}$ and input activations $\mathbf{X}\in\mathbb{R}^{L\times d_{\text{in}}}$, where $d_{\text{in}}$, $d_{\text{out}}$ and $L$ represent weight input dimension, weight output dimension, and input sequence length, respectively. Wanda (Sun et al., 2023) evaluates weight importance through the multiplication of the weight magnitude with the norm of the input feature. Specifically, the importance score for a given weight $\mathbf{W}_{i,j}$ is determined as follows:
@@ 2.2 GLU Variants for Transformer
Gated Linear Units (GLU) (Dauphin et al., 2017) are formed by the element-wise multiplication of two linear projections, with a sigmoid function applied to one projection before the multiplication. Shazeer (2020) suggests an alternative design for the Transformer’s MLP layer that incorporates GLU variants, effectively replacing the conventional first linear transformation and activation function. More formally, we denote the $d_{\text{hidden}}$ as the dimension of the model’s hidden state and $d_{\text{int}}$ as intermediate dimension of the MLP module. For $\mathbf{x}\in\mathbb{R}^{d_{\text{hidden}}}$, $\mathbf{W_{1}}\in\mathbb{R}^{d_{\text{hidden}}\times d_{\text{int}}}$, $\mathbf{W_{2}}\in\mathbb{R}^{d_{\text{hidden}}\times d_{\text{int}}}$, $\mathbf{W_{3}}\in\mathbb{R}^{d_{\text{int}}\times d_{\text{hidden}}}$, and activation function $\sigma(\cdot)$, each MLP layer produces $\mathbf{z}\in\mathbb{R}^{d_{\text{hidden}}}$ by evaluating
@@
We call $\mathbf{W_{1}},\mathbf{W_{2}},\mathbf{W_{3}}$ linear projections as Gate-Proj, Up-Proj, and Down-Proj, respectively. GLU variants that use Swish (Ramachandran et al., 2017) and ReLU (Glorot et al., 2011) activation functions in Eq.(2) are called SwiGLU and ReGLU, respectively. SwiGLU is most widely used in the recent LLMs.
2.3 Revisiting the Effectiveness of Magnitude Pruning
@@
Here, we further investigate the effectiveness of input-balanced magnitude pruning for the MLP module. We divide the SwiGLU-based MLP module into two groups, input projections (Gate-Proj and Up-Proj) and output projection (Down-Proj). We prune those two groups separately and compare Wanda pruning with input-balanced magnitude pruning. Table 1 shows that input-balanced magnitude pruning is surprisingly better than Wanda pruning at higher sparsity ratios for input projections pruning. For the output projection, input-balanced magnitude pruning is consistently worse than Wanda pruning. This suggests the diminished importance of outlier features within input projections compared to the output projection. We analyze the layerwise intermediate activation outlier distribution in Appendix B. However, in the older GeLU-based MLP expansion projection, we observe Wanda can achieve much better performance as shown in Appendix C, suggesting this is a special property of GLU variants.
Up+Gate Projection Down Projection
Sparsity 50% 60% 70% 50% 60% 70%
@@ 2.4 Output-balanced Pruning Limitations
N:M sparsity pattern offers notable speed improvements on recent NVIDIA GPUs (Mishra et al., 2021; Kurtic et al., 2023). It is specified that every group of M consecutive weights must include N zeros. SparseGPT (Frantar & Alistarh, 2023) is not explicitly designed for output-balanced pruning granularity. When converting into an N:M sparsity pattern, SparseGPT forces every M consecutive weight in each row to have exactly N zeros. In such cases, SparseGPT (Frantar & Alistarh, 2023) and Wanda (Sun et al., 2023) unanimously remove the same amount of weights for each output. Such methods often emphasize the significance of individual components within the weight matrix and overlook inter-dependencies to other layers in the network. For MLP input projections pruning, an equal amount of weights corresponding to each intermediate neuron are removed. However, for output projection based on SparseGPT and Wanda pruning metrics, the weights connecting to intermediate neurons with larger activations are more likely to be reserved leading to a structural mismatch.
@@ 3 Dependency-aware Semi-structured Sparsity
In dependency-based structured pruning (Ma et al., 2023; Fang et al., 2023), the initial step is dedicated to recognizing groups of interconnected structures within the model. In terms of SwiGLU-based MLP module pruning, there are three projection matrices, all weights connecting to an identical intermediate neuron are collectively classified into the same dependency group. When pruning weights in $\mathbf{W^{(1)}_{i,:}}$, it is essential to also prune all the weights in $\mathbf{W^{(2)}_{i,:}}$ and $\mathbf{W^{(3)}_{:,i}}$ to maintain the structural consistency of the pruned neural network. In DaSS pruning, instead of aggressively pruning all the weights in less important dependency groups, we incorporate such structural coordination into unstructured pruning. If all weights in $\mathbf{W^{(3)}_{:,i}}$ are emphasized importance through additional importance indicator, such as augmenting corresponding input activations in Eq.(1), then the weights in $\mathbf{W^{(1)}_{i,:}}$ and $\mathbf{W^{(2)}_{i,:}}$ should also be emphasized similar importance.
@@
In dependency-based pruning, we assess the importance of each weight and then aggregate the importance scores within the same group as the group importance score. Consequently, each weight within the same dependency group shares a consistent importance score, ensuring their simultaneous retention or elimination. The importance score of each weight is equal to the group importance score. In DaSS pruning, we take both group importance and weight magnitude into consideration to evaluate the importance of each weight. LLM-pruner (Ma et al., 2023) evaluates the group importance using gradient-based methods. However, computing gradient for LLMs will introduce a significant amount of memory cost and it is less practical for larger models. The existing unstructured pruning methods (Sun et al., 2023; Frantar & Alistarh, 2023) are more efficient than gradient-based methods. By prioritizing the weights linked to outliers in intermediate activations, Wanda outperforms magnitude pruning in Down-Proj pruning. To minimize the impact on these critical outliers, it would be beneficial to also give greater significance to the weights that lead to outlier generation in both Gate-Proj and Up-Proj projections. Thus, we use the norm of intermediate activations $\|\mathbf{y}\|_{2}$ as the group importance indicator. The computation of activation norms is straightforward without introducing more computation and memory overhead than computing gradient. To assign larger importance scores on weights corresponding to more important groups, we evaluate weight importance based on the product of weight magnitude and corresponding group importance score. More specifically, for the Gate-Proj and Up-Proj , the importance score attributed to a specific weight $\mathbf{W}_{i,j}^{k}$ is determined as follows:
@@
In Down-Proj pruning, the pruning metric is the same as done with Wanda (Sun et al., 2023). By augmenting intermediate activations into all three weight importance matrices, DaSS inherently assigns greater emphasis to weights corresponding to intermediate activation outliers, thereby facilitating more nuanced structural coordination among the entire MLP module.
@@
Pruning LLaMA models in finer granularity can improve the performance (Sun et al., 2023). To incorporate intermediate activations into SwiGLU-based MLP pruning, each weight in the same comparison group should correspond to different intermediate activations. In Section 2.3, we show that input-balanced magnitude pruning can still achieve competitive results for Gate-Proj and Up-Proj pruning. In such pruning granularity, we can augment intermediate activations into Gate-Proj and Up-Proj pruning metric. In input-balanced pruning, we remove $s\%$ of the weights linked to each input for a predetermined sparsity ratio of $s\%$ based on weight importance scores. DaSS uses output-balanced sparsity for Down-Proj pruning, which is the same as Wanda.

### v2

# Dependency-Aware Semi-Structured Sparsity: Declining Roles of Outliers in Pruning GLU-based LLMs

@@ Abstract
The rapid growth in the scale of Large Language Models (LLMs) has led to significant computational and memory costs, making model compression techniques such as network pruning increasingly crucial for their efficient deployment. Recent LLMs such as LLaMA2 and Mistral have adopted GLU-based MLP architectures. However, current LLM pruning strategies are primarily based on insights from older LLM architectures, necessitating a reevaluation of these strategies to suit the new architectural characteristics. Contrary to traditional beliefs, we find that outliers play a diminished role in the input projections of GLU-based MLPs. Leveraging this new insight, we propose Dependency-aware Semi-structured Sparsity (DaSS), a novel pruning method for GLU-based LLMs. DaSS balances the flexibility of unstructured pruning and the structural consistency of dependency-based structured pruning by considering both of weight magnitude and corresponding intermediate activation norms in weight pruning metric. Empirical evaluations on the Mistral, Gemma, and LLaMA2 model families demonstrate the consistent effectiveness of DaSS in the prevailing GLU variants.
Dependency-Aware Semi-Structured Sparsity:
Declining Roles of Outliers in Pruning GLU-based LLMs
Zhiyu Guo, Hidetaka Kamigaito, Taro Watanabe Nara Institute of Science and Technology {guo.zhiyu.fy1, kamigaito.h, taro}@is.naist.jp
@@ 1 Introduction
Recent years have witnessed the great success of Large Language Models (LLMs) across various challenging tasks, such as mathematical reasoning, code generation. However, the practical use of these models for inference has faced a major obstacle due to the substantial computational resources they consume. To tackle this, many of the key developments up to now have revolved around weight quantization. It is possible to quantize LLMs down to 4 bits per weight with little impact on accuracy, which aids in memory reduction and speeds up inference Lin et al. (2024). Nonetheless, accuracy recovery becomes problematic when quantizing to around 3 bits per weight with existing methods Dettmers et al. (2024); Egiazarian et al. (2024).
A complementary method is neural network pruning Han et al. (2015b), which can be combined with quantization to further improve the inference efficiency of LLMs Kurtic et al. (2023); Frantar and Alistarh (2023). Pruning can be categorized into two main approaches: unstructured pruning Sun et al. (2024); Frantar and Alistarh (2023), which involves the removal of specific weights, and structured pruning Ma et al. (2023), which entails the removal of complete rows or columns of weights. In contrast to structured pruning, which struggles with performance in LLMs even at low sparsity levels, unstructured pruning methods like SparseGPT Frantar and Alistarh (2023) and Wanda Sun et al. (2024) exhibit promising results without additional retraining, and achieves practical speedup in both CPU and GPU through the recent engineering advancements Agarwalla et al. (2024), they can also benefit in reducing hallucinations of LLMs Chrysostomou et al. (2024).
@@
Figure 1: Illustration of Dependency-aware Semi-structured Sparsity (DaSS). In (a) dependency-based structured pruning Ma et al. (2023), all the weights connecting to the same intermediate neuron are removed or remain simultaneously. In (b) Wanda unstructured pruning Sun et al. (2024), it assigns greater emphasis to the weights corresponding to large input activations. For Gate-Proj and UP-Proj, the same number of weights are removed for each MLP neuron, regardless of whether some neurons have much larger activation norms. For Down-Proj, the weights corresponding to larger activation norms are more likely to be pruned. This can lead to a structural mismatch. In (c), all the weights corresponding to large intermediate activations are more likely to be reserved.
Modern LLMs, such as LLaMA2 Touvron et al. (2023) and Mistral Jiang et al. (2023), have adopted several architectural changes, including the use of GLU (Gated Linear Unit) variants (e.g., SwiGLU) Shazeer (2020) in the MLP modules, and grouped-query attention (GQA) Ainslie et al. (2023). As the GLU-based MLP module accounts for more than 80% of the parameters in LLMs that use GQA  [^1] In LLaMA2-70B, the dimension of key and value is $\frac{1}{8}d$, MLP intermediate dimension is $\frac{7}{2}d$., its pruning emerges as a pivotal factor in determining the overall compression efficacy of LLMs. Notably, prevailing LLM pruning methods such as SparseGPT and Wanda draw inspiration from Llm.int8() Dettmers et al. (2022), which primarily derived its insights from older LLMs like OPT Zhang et al. (2022). However, with the shift in architectural design towards GLU-based MLP modules, it becomes essential to re-evaluate and potentially redefine our LLM pruning strategies to accommodate these new characteristics. Therefore, a timely question arises: which specific properties of GLU-based MLP demand the development of new LLM pruning approaches?
The prevailing belief has been that outlier features are critically important in LLMs Dettmers et al. (2022), and Wanda is a prime example of an outlier-dependent method. Our initial investigation focuses on a more nuanced examination of the effectiveness of Wanda. Sun et al. (2024) suggests that input-balanced magnitude pruning, which prunes the same amount of weights for each input channel, can achieve performance only slightly worse than Wanda. We divide the GLU-based MLP module into two groups: input projections (Gate-Proj and Up-Proj) and output projection (Down-Proj). We compare the effectiveness of input-balanced magnitude pruning and Wanda on these two groups separately. For the input projections, we surprisingly find that input-balanced magnitude pruning performs better than Wanda when sparsity level is higher than 50%. This suggests that outliers play a diminished role in the input projections of GLU-based MLP, contrary to common assumptions. However, for the output projection, Wanda still performs better.
In dependency-aware structured pruning, it’s critical to consider that pruned parameters have dependencies with other parameters, owing to their interconnected nature Ma et al. (2023); Fang et al. (2023). For GLU-based MLP pruning, all the weights connected to the same intermediate neuron should be preserved or pruned simultaneously.
@@
Leveraging our newfound insight, and drawing inspiration from dependency-based structured pruning method, we introduce a new LLM pruning paradigm, namely, Dependency-aware Semi-structured Sparsity (DaSS). This approach is specifically designed to navigate the middle ground between the flexibility of unstructured pruning and the structural consistency of dependency-based structured pruning. We present a new MLP pruning metric that assesses each weight’s importance based on the product of its magnitude and the norm of the corresponding MLP intermediate activations. Our proposed DaSS method, illustrated in Figure 1(c), embodies a semi-structured pattern that retains a degree of the adaptability inherent in unstructured pruning while incorporating the dependency-aware aspect of structured pruning. This balance allows for more precise pruning.
We perform extensive experiments on LLaMA2  Touvron et al. (2023), Gemma  Team et al. (2024), and Mistral  Jiang et al. (2023) to evaluate DaSS across various tasks from language modeling, 5 commonsense reasoning tasks. In achieving hardware-friendly N:M sparsity patterns, DaSS consistently excels beyond the existing LLM pruning methods SparseGPT and Wanda, while maintaining the computational efficiency akin to Wanda. Moreover, DaSS demonstrates consistent effectiveness in all the prevalent GLU variants, including SwiGLU, GeGLU and ReGLU. Impressively, even without weight updates used in SparseGPT, we observe significant improvements achieved by DaSS over Wanda with LLaMA2-70B in 5 commonsense reasoning tasks, with mean accuracy improvement of 3.63 and 12.24 at sparsity levels of 70% and 80%, respectively. Note that the MLP input projections account for around 50% parameters in GQA-based LLMs, DaSS prunes them without using input activation, challenging the traditional wisdom that outliers play a pivotal role in LLM compression. We hope our fresh insights can also motivate more nuanced GLU-specific LLM quantization strategies.
2 Related Work
Pruning LLM.
Neural network pruning in LLM can be broadly categorized into two groups: structured pruning Ma et al. (2023); Xia et al. (2023a) and unstructured pruning Frantar and Alistarh (2023); Sun et al. (2024). Ma et al. (2023) proposes a dependency detection algorithm to detect and prune non-critical grouped structures. Although structured pruning can have better hardware efficiency, the accuracy drops a lot even at a low compression rate, they typically require re-trianing to recover accuracy. Unstructured pruning can yield a higher compression rate and achieve acceleration on both CPUs and GPUs through the recent software and handware advancement Agarwalla et al. (2024). SparseGPT Frantar and Alistarh (2023) leverages the Hessian inverse for pruning and reduces reconstruction error of dense and sparse weights by subsequent weight updates. Wanda Sun et al. (2024) employs an efficient method that augments input activations into weight magnitudes, and matches the performance of SparseGPT at medium sparsity. OWL Yin et al. (2024) explores non-uniform layer-wise sparsity for pruning LLMs, however, it is less practical to be converted into N:M sparsity. Our work incorporates dependency information into unstructured pruning, achieving a novel pruning paradigm.
Outlier-dependent LLM Compression.
Outlier features, defined as features with magnitudes substantially larger than others, are a notable characteristic of LLMs Dettmers et al. (2022). Despite making up only a small fraction of all feature dimensions, these outliers play a critical role in attention and predictive performance. Such observation has motivated the development of LLM-specific quantization methods Dettmers et al. (2022); Xiao et al. (2023); Lin et al. (2024); Ashkboos et al. (2023) to handle outliers more effectively. Wanda Sun et al. (2024) and OWL Yin et al. (2024) broaden these insights, revealing that outlier features are significant in deciding weight importance when pruning LLMs. Our analysis diverges from conventional wisdom by demonstrating that, in the context of GLU-based MLP input projections, the influence of outlier features is not as pronounced as previously assumed, prompting a reevaluation of their role in LLM compression strategies.
@@ 3.1 Wanda Pruning Method
In the context of LLM pruning, we denote a linear layer weight matrix $\mathbf{W}\in\mathbb{R}^{d_{\text{out}}\times d_{\text{in}}}$ and input activations $\mathbf{X}\in\mathbb{R}^{L\times d_{\text{in}}}$, where $d_{\text{in}}$, $d_{\text{out}}$ and $L$ represent weight input dimension, weight output dimension, and input sequence length, respectively. Wanda Sun et al. (2024) evaluates weight importance through the multiplication of the weight magnitude with the norm of the input feature. Specifically, the importance score for a given weight $\mathbf{W}_{i,j}$ is determined as follows:
@@ 3.2 GLU Variants for Transformer
Gated Linear Units (GLU) Dauphin et al. (2017) are formed by the element-wise multiplication of two linear projections, with a sigmoid function applied to one projection before the multiplication. Shazeer (2020) suggests an alternative design for the Transformer’s MLP layer that incorporates GLU variants, effectively replacing the conventional first linear transformation and activation function. More formally, we denote the $d_{\text{hidden}}$ as the dimension of the model’s hidden state and $d_{\text{int}}$ as intermediate dimension of the MLP module. For $\mathbf{x}\in\mathbb{R}^{d_{\text{hidden}}}$, $\mathbf{W_{1}}\in\mathbb{R}^{d_{\text{hidden}}\times d_{\text{int}}}$, $\mathbf{W_{2}}\in\mathbb{R}^{d_{\text{hidden}}\times d_{\text{int}}}$, $\mathbf{W_{3}}\in\mathbb{R}^{d_{\text{int}}\times d_{\text{hidden}}}$, and activation function $\sigma(\cdot)$, each MLP layer produces $\mathbf{z}\in\mathbb{R}^{d_{\text{hidden}}}$ by evaluating
@@
We call $\mathbf{W_{1}},\mathbf{W_{2}},\mathbf{W_{3}}$ linear projections as Gate-Proj, Up-Proj, and Down-Proj, respectively. GLU variants that use Swish Ramachandran et al. (2017), GeLU Hendrycks and Gimpel (2016), and ReLU Glorot et al. (2011) activation functions in Eq.(2) are called SwiGLU, GeGLU, and ReGLU, respectively. SwiGLU is most widely used in the recent LLMs.
3.3 Revisiting the Effectiveness of Wanda
@@
Here, we further investigate the effectiveness of input-balanced magnitude pruning for the MLP module. We divide the GLU-based MLP module into two groups, input projections (Gate-Proj and Up-Proj) and output projection (Down-Proj). We prune those two groups separately and compare Wanda pruning with input-balanced magnitude pruning. Table 1 shows that input-balanced magnitude pruning is surprisingly better than Wanda pruning at higher sparsity ratios for input projections pruning. For the output projection, input-balanced magnitude pruning is consistently worse than Wanda pruning. This suggests the diminished importance of outlier features within input projections compared to the output projection. However, in the older GeLU-based MLP expansion projection, we observe Wanda can achieve much better performance as shown in Appendix A.3, suggesting this is a special property of GLU variants.
Method Up+Gate Projection Down Projection
50% 60% 70% 50% 60% 70%
@@ 3.4 Output-balanced Pruning Limitations
N:M sparsity pattern offers notable speed improvements on recent NVIDIA GPUs Mishra et al. (2021); Kurtic et al. (2023). It is specified that every group of M consecutive weights must include N zeros. SparseGPT Frantar and Alistarh (2023) is not explicitly designed for output-balanced pruning granularity. When converting into an N:M sparsity pattern, SparseGPT forces every M consecutive weight in each row to have exactly N zeros. In such cases, SparseGPT Frantar and Alistarh (2023) and Wanda Sun et al. (2024) unanimously remove the same amount of weights for each output. Such methods often emphasize the significance of individual components within the weight matrix and overlook inter-dependencies to other layers in the network. For MLP input projections pruning, an equal amount of weights corresponding to each intermediate neuron are removed. However, for output projection based on SparseGPT and Wanda pruning metrics, the weights connecting to intermediate neurons with larger activations are more likely to be reserved leading to a structural mismatch.
@@ 4 Dependency-aware Semi-structured Sparsity
In dependency-based structured pruning Ma et al. (2023); Fang et al. (2023), the initial step is dedicated to recognizing groups of interconnected structures within the model. In terms of GLU-based MLP module pruning, there are three projection matrices, all weights connecting to an identical intermediate neuron are collectively classified into the same dependency group. When pruning weights in $\mathbf{W^{(1)}_{i,:}}$, it is essential to also prune all the weights in $\mathbf{W^{(2)}_{i,:}}$ and $\mathbf{W^{(3)}_{:,i}}$ to maintain the structural consistency of the pruned neural network. In DaSS pruning, instead of aggressively pruning all the weights in less important dependency groups, we incorporate such structural coordination into unstructured pruning. If all weights in $\mathbf{W^{(3)}_{:,i}}$ are emphasized importance through additional importance indicator, such as augmenting corresponding input activations in Eq.(1), then the weights in $\mathbf{W^{(1)}_{i,:}}$ and $\mathbf{W^{(2)}_{i,:}}$ should also be emphasized similar importance.
@@
In dependency-based pruning, we assess the importance of each weight and then aggregate the importance scores within the same group as the group importance score. Consequently, each weight within the same dependency group shares a consistent importance score, ensuring their simultaneous retention or elimination. The importance score of each weight is equal to the group importance score. In DaSS pruning, we take both group importance and weight magnitude into consideration to evaluate the importance of each weight. LLM-pruner Ma et al. (2023) evaluates the group importance using gradient-based methods. However, computing gradient for LLMs will introduce a significant amount of memory cost and it is less practical for larger models. The existing unstructured pruning methods Sun et al. (2024); Frantar and Alistarh (2023) are more efficient than gradient-based methods. By prioritizing the weights linked to outliers in intermediate activations, Wanda outperforms magnitude pruning in Down-Proj pruning. To minimize the impact on these critical outliers, it would be beneficial to also give greater significance to the weights that lead to outlier generation in both Gate-Proj and Up-Proj projections. Thus, we use the norm of intermediate activations $\|\mathbf{y}\|_{2}$ as the group importance indicator. The computation of activation norms is straightforward without introducing more computation and memory overhead than computing gradient. To assign larger importance scores on weights corresponding to more important groups, we evaluate weight importance based on the product of weight magnitude and corresponding group importance score. More specifically, for the Gate-Proj and Up-Proj , the importance score attributed to a specific weight $\mathbf{W}_{i,j}^{k}$ is determined as follows:
@@
In Down-Proj pruning, the pruning metric is the same with Wanda Sun et al. (2024). By augmenting intermediate activations into all three weight importance matrices, DaSS inherently assigns greater emphasis to weights corresponding to intermediate activation outliers, thereby facilitating more nuanced structural coordination among the entire MLP module.
@@
Pruning LLaMA models in finer granularity can improve the performance Sun et al. (2024). To incorporate intermediate activations into SwiGLU-based MLP pruning, each weight in the same comparison group should correspond to different intermediate activations. In Section 2.3, we show that input-balanced magnitude pruning can still achieve competitive results for Gate-Proj and Up-Proj pruning. In such pruning granularity, we can augment intermediate activations into Gate-Proj and Up-Proj pruning metric. In input-balanced pruning, we remove $s\%$ of the weights linked to each input for a predetermined sparsity ratio of $s\%$ based on weight importance scores. DaSS uses output-balanced sparsity for Down-Proj pruning, which is the same as Wanda.